# Nonflat bands and chiral symmetry in magic-angle twisted bilayer graphene

### v1

@@ IV Symmetry of MATBG
We will argue now that the matrix element ${{H}}_{\text{int},\alpha^{\prime}\beta^{\prime}\gamma^{\prime}\delta^{\prime}}$ between modes $\eta^{\prime}=C_{2z}\tilde{P}|\eta\rangle$ ($\eta=\alpha,\beta,\gamma,\delta$) is approximately equal to ${{H}}_{\text{int},\alpha\beta\gamma\delta}$. For that, we may establish a one-to-one correspondence between atoms $\bm{r}$ and $\bm{s}$, with $\bm{s}$ belonging to the opposite sublattice and layer to $\bm{r}$ and with approximately the same in-plane coordinates. Then, we can replace the sums over $\bm{r},\bm{r^{\prime}}$ in Eq. (S18) by sums over $\bm{s},\bm{s^{\prime}}$ and conclude that, in the $U(2)_{K}\times U(2)_{K^{\prime}}$ limit where $\tau_{\alpha}=\tau_{\delta}$, $\tau_{\beta}=\tau_{\gamma}$ and the valley phases cancel, ${{H}}_{\text{int},\alpha\beta\gamma\delta}={{H}}_{\text{int},\alpha^{\prime}\beta^{\prime}\gamma^{\prime}\delta^{\prime}}$ to a very good approximation. This property hints to the existence of a continuous symmetry generated by $C_{2z}\tilde{P}$. Indeed, in Appendix D of Ref. [102] we showed that $[{H}_{\text{int}},C_{2z}\tilde{P}]=0$ (assuming $U(2)_{K}\times U(2)_{K^{\prime}}$ and up to effects from the discrete lattice), also we provide more details on the implementation of $C_{2z}\tilde{P}$ in the atomistic model in Appendix C of Ref. [102].
@@
With regards to the local potential, in the case of perfect particle-hole symmetry, and neglecting intervalley scatterings, we have $C_{2z}{P}\rho_{\bm{r}s}C_{2z}{P}=\rho_{\bm{s}s}$. Then, provided $\rho_{0\bm{r}}=\rho_{0\bm{s}}$, the potential respects the $C_{2z}P$ symmetry. We find that $\rho_{0\bm{r}}\approx\rho_{0\bm{s}}$ to a good approximation, as shown in Fig. S2(b); moreover we also find that $\rho_{0\bm{r}}\approx\frac{1}{2}\sum_{\bm{k}n}^{\prime}|\langle\bm{r}|\bm{k}n\rangle|^{2}$. These properties of $\rho_{0\bm{r}}$ can be partly understood from the constrains imposed by the crystallographic and valley symmetries and the properties of the normal state under $C_{2z}P$ and $\sigma_{z}$, see Eq. (IV) (where we can replace $\mathcal{H}_{0}$ by the normal state, $H_{\text{MF}}$). Notice also that in $\pi(\bm{r})$ the density is convoluted with the Coulomb potential, which tends to smoothen the inhomogeneities making the relation $\pi(\bm{r})\approx\pi(\bm{s})$ tighter. Particle-hole asymmetry, $C_{2z}\overline{P}\neq C_{2z}P$, translates to some amount of symmetry breaking.
@@
$\big{[}C_{2z}P,{\mathcal{H}}_{\text{int}}\big{]}\approx 0,$ (S23)
and the strength of symmetry breaking can be assessed by the singular values of $C_{2z}\overline{P}$, displayed in Fig. S2(a) and Table (S2).
@@
$\tilde{R}=C_{2z}\tilde{P}\tilde{\sigma}_{z}\tau_{z}=\sigma_{y}\tau_{x}\mu_{y},$ (S24)
@@
$\big{[}R,{\mathcal{H}}_{\text{int}}\big{]}\approx 0,$ (S25)
@@
Once identified the ’hidden’ symmetries, we can obtain the full symmetry group by commuting the $U(2)_{K}\times U(2)_{K^{\prime}}$ generators of Eq. (S20) with $C_{2z}P$, $R$ and/or $\sigma_{z}\tau_{z}$. For simplicity, let us consider the spinless system, in which $U(2)_{K}\times U(2)_{K^{\prime}}$ turns to $U(1)_{\text{C}}\times U(1)_{\text{V}}$, generated by the identity, $\mathbb{1}$, and $\tau_{z}$. Let us also work in the Chern basis of the $n_{B}=4$ model for concreteness. The symmetry generators are written as $\mathcal{G}=\sum_{\bm{k}}c^{\dagger}_{\bm{k}}Gc_{\bm{k}}$ with $c^{\dagger}_{\bm{k}}=\big{(}c^{\dagger}_{\bm{k}\alpha}\big{)}$ a vector and $G=[G]_{\alpha\beta}$ a matrix in the Chern basis; let us work with just the $G$ matrices for neatness.
@@
The same procedure can be followed straightforwardly including the spin; the $U(2)$ groups are enlarged to $U(4)$ and the $U(2)\times U(2)$ to $U(4)\times U(4)$. The different limits and their symmetry groups are listed in Table (S1). The case with $U(2)_{K}\times U(2)_{K^{\prime}}$ and $C_{2z}P$ is dubbed ’flat’ limit and the symmetry $U(4)_{\text{flat}}$ (the dispersion $\mathcal{H}_{0}$ is prone to breaking this symmetry, as we discuss below). On the other hand, with $U(2)_{K}\times U(2)_{K^{\prime}}$ and $R$ we are in the ’chiral-nonflat’ limit and the symmetry is denoted $U(4)_{\text{chiral}}$. Note that in the main text we use the notations $C_{2z}P=\mathcal{G}_{\text{flat}}$ and $R=\mathcal{G}_{\text{chiral}}$.
@@
Table S1: Different limits and their symmetry groups in the Chern basis of $n_{B}=4$ MATBG. The explicit symmetry breaking hierarchy is $U(4)\times U(4)\to U(4)_{\text{chiral}}\to U(2)_{K}\times U(2)_{K^{\prime}}\to SU(2)_{\text{spin}}\times U(1)_{\text{V}}\times U(1)_{\text{C}}$ as we have discussed in this work.
@@
We see that only in the chiral limit the $\tilde{R}$ symmetry remains. As a matter of fact, the relations $\{H_{0},C_{2z}\tilde{P}\}=0$ and (approximately) $\{H_{0},\tilde{\sigma}_{z}\tau_{z}\}=0$ imply that the symmetry is broken by energy scales as large as the full bandwidth of the theory. Projecting to the flat bands effectively reduces the bandwidth and in strong coupling, $H_{\text{int}}\gg H_{0}\approx 0$, we have $[C_{2z}\tilde{P},H_{0}]=[C_{2z}\tilde{P},H_{\text{int}}]=0$ and additionally $[\tilde{\sigma}_{z}\tau_{z},H_{0}]=[\tilde{\sigma}_{z}\tau_{z},H_{\text{int}}]=0$ in the chiral limit. Nonetheless, a proper many-body projection to the flat bands will presumably induce new components in $H_{0}$ as we have shown in this work, disallowing the approximation $H_{0}\approx 0$.
@@
${h}(\bm{k})$ $={h}_{00}(\bm{k})+{h}_{z0}(\bm{k})\tau_{z}+{h}_{x0}(\bm{k})\sigma_{x}+{h}_{yz}(\bm{k})\sigma_{y}\tau_{z}+{h}_{y0}(\bm{k})\sigma_{y}+{h}_{xz}(\bm{k})\sigma_{x}\tau_{z},$ (S26)
The strength of the different components is summarized in Table (S2). The dominant term is $h_{x0}(\bm{k})\sigma_{x}+h_{yz}(\bm{k})\sigma_{y}\tau_{z}$, with ${h}_{00}(\bm{k})$ also sizable at the $\Gamma_{M}$ point. In fact, $h_{x0}(\bm{k})$ and $h_{yz}(\bm{k})$ are the main sources of symmetry breaking in $n_{B}=4$ MATBG, as we discuss in the main text. We have then the relations (neglecting $h_{00}$)
$\big{\{}C_{2z}P,\mathcal{H}_{0}\big{\}}\approx 0,\quad\quad\quad\big{\{}\sigma_{z}\tau_{z},\mathcal{H}_{0}\big{\}}\approx 0,\quad\quad\quad\big{[}R,\mathcal{H}_{0}\big{]}\approx 0.$ (S27)
@@
Then, $\mathcal{H}_{0}$ breaks the $C_{2z}P$ and $\sigma_{z}\tau_{z}$ symmetry. All the more, in theories with larger bandwidth, say $n_{B}=20$, and assuming that the relations of Eq. (S27) are approximately satisfied, the $C_{2z}P$ and $\sigma_{z}\tau_{z}$ symmetries are more seriously broken. In this respect, the authors of Ref. [68] discovered a chiral operator having a simple form in the THF basis, denoted $\tilde{S}$, with $[\tilde{S},H_{\text{int}}]\approx 0$ and $\{\tilde{S},H_{0}\}\approx 0$; the non-anticommuting terms being of the order of the flat bandwidth. This ’third chiral’ operator can be combined with the particle-hole operator resulting in $[C_{2z}\tilde{P}\tilde{S},H_{0}]\approx 0$ and $[C_{2z}\tilde{P}\tilde{S},H_{\text{int}}]\approx 0$. It was shown that $\tilde{S}$ is consistent with the identity in (most of) the flat bands, resulting in that $C_{2z}\tilde{P}\tilde{S}$ approximately anticommutes with the flat bands and approximately commutes with the higher bands of $H_{0}$. This is in contrast to $C_{2z}\tilde{P}$ that anticommutes with all the bands. Similarly, the operator $\tilde{\sigma}_{z}\tau_{z}\tilde{S}$ approximately commutes with $H_{\text{int}}$ and with the high-energy bands of $H_{0}$ (even though the relation $\{\tilde{\sigma}_{z},H_{0}\}\approx 0$ is expected to become worse on the higher bands), and approximately anticommutes with the flat bands.
@@
Table S2: Relevant figures for the breaking of the $U(4)\times U(4)$ symmetry in the $n_{B}=4$ system. Singular values of $C_{2z}\overline{P}$ different from $1$ break the flat symmetry of $\mathcal{H}_{\text{int}}$, and eigenvalues of $\overline{\sigma}_{z}$ different from $1$ break the chiral symmetry of $\mathcal{H}_{\text{int}}$. The components $h_{0z}(\bm{k})$, $h_{x0}(\bm{k})$ and $h_{yz}(\bm{k})$ break the flat symmetry of $\mathcal{H}_{0}$, and the components $h_{0z}(\bm{k})$, $h_{y0}(\bm{k})$ and $h_{xz}(\bm{k})$ break the chiral symmetry of $\mathcal{H}_{0}$. avBZ denotes the average over the Brillouin zone.
V Topological phase transition in the graphene subtraction scheme
@@
${H}=H_{\text{TB}}+{H}_{\text{int}}$ (S28)
$H_{\text{TB}}=\sum_{\bm{r}\bm{r^{\prime}}s}t(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}$ (S29)
@@
$\hskip 29.87547pt-\sum_{\bm{r}s}U\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{\text{gr}}\ c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\text{constant},$ (S30)
@@
$-\sum_{\bm{r}\neq\bm{r^{\prime}},s}V(\bm{r}-\bm{r})\bigg{(}\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{0}-\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{\text{gr}}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\sum_{\bm{r}s}U\bigg{(}\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{0}-\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{\text{gr}}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\text{constant}.$ (S31)
@@
Contrary to this expectation, our calculations with the graphene subtraction also display significant band widening, as depicted in Fig. S3. All the more, for some value of $\epsilon_{r}$ between $50$ and $10$ we find a topological phase transition moving the one-dimensional (two-dimensional) representations (per valley) at $\Gamma_{M}$ from the flat (remote) bands to the remote (flat) bands; the spectral weight redistribution in the graphene subtraction scheme is severe. Notice that the flat bands become entangled with the higher bands making the $n_{B}=4$ projector ill-defined.
@@
Figure S3: Band structures of the symmetric state in the graphene subtraction scheme for $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0,5$ eV, compared to the non-interacting (bare) bands. The system underwent a topological phase transition between $\epsilon_{r}=50$ and $10$ and the flat bands become doubly (per valley) degenerate at $\Gamma_{M}$.
VI Additional results: many-body projection

### v3

@@ Appendix D Symmetry of MATBG
We will argue now that the matrix element ${{H}}_{\text{int},\alpha^{\prime}\beta^{\prime}\gamma^{\prime}\delta^{\prime}}$ between modes $\eta^{\prime}=C_{2z}\tilde{P}|\eta\rangle$ ($\eta=\alpha,\beta,\gamma,\delta$) is approximately equal to ${{H}}_{\text{int},\alpha\beta\gamma\delta}$. For that, we may establish a one-to-one correspondence between atoms $\bm{r}$ and $\bm{s}$, with $\bm{s}$ belonging to the opposite sublattice and layer to $\bm{r}$ and with approximately the same in-plane coordinates. Then, we can replace the sums over $\bm{r},\bm{r^{\prime}}$ in Eq. (22) by sums over $\bm{s},\bm{s^{\prime}}$ and conclude that, in the $U(2)_{K}\times U(2)_{K^{\prime}}$ limit where $\tau_{\alpha}=\tau_{\delta}$, $\tau_{\beta}=\tau_{\gamma}$ and the valley phases cancel, ${{H}}_{\text{int},\alpha\beta\gamma\delta}={{H}}_{\text{int},\alpha^{\prime}\beta^{\prime}\gamma^{\prime}\delta^{\prime}}$ to a very good approximation. This property hints to the existence of a continuous symmetry generated by $C_{2z}\tilde{P}$. Indeed, in Appendix D of Ref. [106] we showed that $[{H}_{\text{int}},C_{2z}\tilde{P}]=0$ (assuming $U(2)_{K}\times U(2)_{K^{\prime}}$ and up to effects from the discrete lattice), also we provide more details on the implementation of $C_{2z}\tilde{P}$ in the atomistic model in Appendix C of Ref. [106].
@@
With regards to the local potential, in the case of perfect particle-hole symmetry, and neglecting intervalley scatterings, we have $C_{2z}{P}\rho_{\bm{r}s}C_{2z}{P}=\rho_{\bm{s}s}$. Then, provided $\rho_{0\bm{r}}=\rho_{0\bm{s}}$, the potential respects the $C_{2z}P$ symmetry. We find that $\rho_{0\bm{r}}\approx\rho_{0\bm{s}}$ to a good approximation, as shown in Fig. D2(b); moreover we also find that $\rho_{0\bm{r}}\approx\frac{1}{2}\sum_{\bm{k}n}^{\prime}|\langle\bm{r}|\bm{k}n\rangle|^{2}$. These properties of $\rho_{0\bm{r}}$ can be partly understood from the constrains imposed by the crystallographic and valley symmetries and the properties of the normal state under $C_{2z}P$ and $\sigma_{z}$, see Eq. (D) (where we can replace $\mathcal{H}_{0}$ by the normal state, $H_{\text{MF}}$). Notice also that in $\pi(\bm{r})$ the density is convolved with the Coulomb potential, which tends to smoothen the inhomogeneities making the relation $\pi(\bm{r})\approx\pi(\bm{s})$ tighter. Particle-hole asymmetry, $C_{2z}\overline{P}\neq C_{2z}P$, translates to some amount of symmetry breaking.
@@
$\big{[}C_{2z}P,{\mathcal{H}}_{\text{int}}\big{]}\approx 0,$ (27)
and the strength of symmetry breaking can be assessed by the singular values of $C_{2z}\overline{P}$, displayed in Fig. D2(a) and Table (D2).
@@
$\tilde{R}=C_{2z}\tilde{P}\tilde{\sigma}_{z}\tau_{z}=\sigma_{y}\tau_{x}\mu_{y},$ (28)
@@
$\big{[}R,{\mathcal{H}}_{\text{int}}\big{]}\approx 0,$ (29)
@@
Once identified the ’hidden’ symmetries, we can obtain the full symmetry group by commuting the $U(2)_{K}\times U(2)_{K^{\prime}}$ generators of Eq. (24) with $C_{2z}P$, $R$ and/or $\sigma_{z}\tau_{z}$. For simplicity, let us consider the spinless system, in which $U(2)_{K}\times U(2)_{K^{\prime}}$ turns to $U(1)_{\text{C}}\times U(1)_{\text{V}}$, generated by the identity, $\mathbb{1}$, and $\tau_{z}$. Let us also work in the Chern basis of the $n_{B}=4$ model for concreteness. The symmetry generators are written as $\mathcal{G}=\sum_{\bm{k}}c^{\dagger}_{\bm{k}}Gc_{\bm{k}}$ with $c^{\dagger}_{\bm{k}}=\big{(}c^{\dagger}_{\bm{k}\alpha}\big{)}$ a vector and $G=[G]_{\alpha\beta}$ a matrix in the Chern basis; let us work with just the $G$ matrices for neatness.
@@
The same procedure can be followed straightforwardly including the spin; the $U(2)$ groups are enlarged to $U(4)$ and the $U(2)\times U(2)$ to $U(4)\times U(4)$. The different limits and their symmetry groups are listed in Table (D1). The case with $U(2)_{K}\times U(2)_{K^{\prime}}$ and $C_{2z}P$ is dubbed ’flat’ limit and the symmetry $U(4)_{\text{flat}}$ (the dispersion $\mathcal{H}_{0}$ is prone to breaking this symmetry, as we discuss below). On the other hand, with $U(2)_{K}\times U(2)_{K^{\prime}}$ and $R$ we are in the ’chiral-nonflat’ limit and the symmetry is denoted $U(4)_{\text{chiral}}$. Note that in the main text we use the notations $C_{2z}P=\mathcal{G}_{\text{flat}}$ and $R=\mathcal{G}_{\text{chiral}}$.
@@
Table D1: Different limits and their symmetry groups in the Chern basis of $n_{B}=4$ MATBG. The explicit symmetry breaking hierarchy is $U(4)\times U(4)\to U(4)_{\text{chiral}}\to U(2)_{K}\times U(2)_{K^{\prime}}\to SU(2)_{\text{spin}}\times U(1)_{\text{V}}\times U(1)_{\text{C}}$ as we have discussed in this work.
@@
We see that only in the chiral limit the $\tilde{R}$ symmetry remains. As a matter of fact, the relations $\{H_{0},C_{2z}\tilde{P}\}=0$ and (approximately) $\{H_{0},\tilde{\sigma}_{z}\tau_{z}\}=0$ imply that the symmetry is broken by energy scales as large as the full bandwidth of the theory. Projecting to the flat bands effectively reduces the bandwidth and in strong coupling, $H_{\text{int}}\gg H_{0}\approx 0$, we have $[C_{2z}\tilde{P},H_{0}]=[C_{2z}\tilde{P},H_{\text{int}}]=0$ and additionally $[\tilde{\sigma}_{z}\tau_{z},H_{0}]=[\tilde{\sigma}_{z}\tau_{z},H_{\text{int}}]=0$ in the chiral limit. Nonetheless, a proper many-body projection to the flat bands will presumably induce new components that widen the bands of $H_{0}$ at the bare magic angle as we have shown in this work, disallowing the approximation $H_{0}\approx 0$ (at the renormalized magic angle the induce terms will tend to cancel the bare $H_{0}$ thus recovering the strong coupling limit).
@@
${h}(\bm{k})$ $={h}_{00}(\bm{k})+{h}_{z0}(\bm{k})\tau_{z}+{h}_{x0}(\bm{k})\sigma_{x}+{h}_{yz}(\bm{k})\sigma_{y}\tau_{z}+{h}_{y0}(\bm{k})\sigma_{y}+{h}_{xz}(\bm{k})\sigma_{x}\tau_{z},$ (30)
The strength of the different components is summarized in Table (D2). The dominant term is $h_{x0}(\bm{k})\sigma_{x}+h_{yz}(\bm{k})\sigma_{y}\tau_{z}$, with ${h}_{00}(\bm{k})$ also sizable at the $\Gamma_{M}$ point. In fact, $h_{x0}(\bm{k})$ and $h_{yz}(\bm{k})$ are the main sources of symmetry breaking as we discuss in the main text. We have then the relations (neglecting $h_{00}$)
$\big{\{}C_{2z}P,\mathcal{H}_{0}\big{\}}\approx 0,\quad\quad\quad\big{\{}\sigma_{z}\tau_{z},\mathcal{H}_{0}\big{\}}\approx 0,\quad\quad\quad\big{[}R,\mathcal{H}_{0}\big{]}\approx 0.$ (31)
@@
Then, $\mathcal{H}_{0}$ breaks the $C_{2z}P$ and $\sigma_{z}\tau_{z}$ symmetry. All the more, in theories with larger bandwidth, say $n_{B}=20$, and assuming that the relations of Eq. (31) are approximately satisfied, the $C_{2z}P$ and $\sigma_{z}\tau_{z}$ symmetries are more seriously broken. In this respect, the authors of Ref. [70] discovered a chiral operator having a simple form in the THF basis, denoted $\tilde{S}$, with $[\tilde{S},H_{\text{int}}]\approx 0$ and $\{\tilde{S},H_{0}\}\approx 0$; the non-anticommuting terms being of the order of the flat bandwidth. This ’third chiral’ operator can be combined with the particle-hole operator resulting in $[C_{2z}\tilde{P}\tilde{S},H_{0}]\approx 0$ and $[C_{2z}\tilde{P}\tilde{S},H_{\text{int}}]\approx 0$. It was shown that $\tilde{S}$ is consistent with the identity in (most of) the flat bands, resulting in that $C_{2z}\tilde{P}\tilde{S}$ approximately anticommutes with the flat bands and approximately commutes with the higher bands of $H_{0}$. This is in contrast to $C_{2z}\tilde{P}$ that anticommutes with all the bands. Similarly, the operator $\tilde{\sigma}_{z}\tau_{z}\tilde{S}$ approximately commutes with $H_{\text{int}}$ and with the high-energy bands of $H_{0}$ (even though the relation $\{\tilde{\sigma}_{z},H_{0}\}\approx 0$ is expected to become worse on the higher bands), and approximately anticommutes with the flat bands.
@@
Table D2: Relevant figures for the breaking of the $U(4)\times U(4)$ symmetry in the $n_{B}=4$ system. Singular values of $C_{2z}\overline{P}$ different from $1$ break the flat symmetry of $\mathcal{H}_{\text{int}}$, and eigenvalues of $\overline{\sigma}_{z}$ different from $1$ break the chiral symmetry of $\mathcal{H}_{\text{int}}$. The components $h_{0z}(\bm{k})$, $h_{x0}(\bm{k})$ and $h_{yz}(\bm{k})$ break the flat symmetry of $\mathcal{H}_{0}$, and the components $h_{0z}(\bm{k})$, $h_{y0}(\bm{k})$ and $h_{xz}(\bm{k})$ break the chiral symmetry of $\mathcal{H}_{0}$. avBZ denotes the average over the Brillouin zone.
Appendix E Topological phase transition in the graphene subtraction scheme
@@
${H}=H_{\text{TB}}+{H}_{\text{int}}$ (32)
$H_{\text{TB}}=\sum_{\bm{r}\bm{r^{\prime}}s}t(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}$ (33)
@@
$\hskip 29.87547pt-\sum_{\bm{r}s}U\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{\text{gr}}\ c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\text{constant},$ (34)
@@
$-\sum_{\bm{r}\neq\bm{r^{\prime}},s}V(\bm{r}-\bm{r})\bigg{(}\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{0}-\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{\text{gr}}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\sum_{\bm{r}s}U\bigg{(}\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{0}-\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{\text{gr}}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\text{constant}.$ (35)
@@
Contrary to this expectation, our calculations with the graphene subtraction also display significant band widening, as depicted in Fig. E3. All the more, for some value of $\epsilon_{r}$ between $50$ and $10$ we find a topological phase transition moving the one-dimensional (two-dimensional) representations (per valley) at $\Gamma_{M}$ from the flat (remote) bands to the remote (flat) bands. Notice that the flat bands become entangled with the higher bands making the $n_{B}=4$ projector ill-defined.
@@
Figure E3: Band structures of the symmetric state in the graphene subtraction scheme for $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0.5$ eV, compared to the non-interacting (bare) bands. The system underwent a topological phase transition between $\epsilon_{r}=50$ and $10$ and the flat bands become doubly (per valley) degenerate at $\Gamma_{M}$. The inset shows the Brillouin zone and the momentum lines plotted.
Appendix F Additional information: many-body projection
The many-body projection introduced in Appendix B relies on the band projector onto excitations of the central $n_{B}$ bands, Eq. 10. This projector is well defined if the set of central bands is gapped from the remaining bands, and it can be approximated numerically via a discretization of the Brillouin zone. In MATBG we can consider a weaker condition, which is that the central $n_{B}/2$ bands are gapped from the remaining bands for each valley, because the two valleys are completely decoupled
For the $n_{B}=4$ theory, the flat bands are clearly isolated from the remote bands, with gaps of some tens of meV. Let us focus then on $n_{B}=20$. In Fig. F4(a) we plot the band structures of the symmetric state along the high-symmetry $K_{M}\Gamma_{M}M_{M}K_{M}$ line; the lowest $20$ bands are shown in blue. Clearly, the set of the first $20$ bands is gapped from the remaining bands at the high symmetry points $\Gamma_{M}$, $K_{M}$ and $M_{M}$. Moreover, the symmetry representations are consistent with those of a topologically trivial system with ten orbitals per valley [54, 62].
In Fig. F4(b) we show the gaps on the electron and hole sides on the Brillouin zone. The values correspond to the valley $K$ bands, i.e. bands with positive $\tilde{\tau}_{z}$ expectation value (Eq. 18); valley $K^{\prime}$ is obtained by a $180^{\circ}$ rotation. On the left column we consider the $6\times 6$ discretization used in our self-consistent algorithm. As anticipated by the analysis of the high-symmetry points, the gaps remain finite for all $\bm{k}$, hence the projector to the central 10 bands per valley (equivalently, the central $20$ bands) in the $6\times 6$ system is unambiguously defined and respects the symmetries of the system.
In an infinitely fine momentum grid band touchings appear at generic momenta, near the $\Gamma_{M}M_{M}$ line (or a symmetry-related line), and the band projector becomes ill-defined; these touchings are noticeable on the right column of Fig. F4(b) for a $96\times 96$ grid. One can formally circumvent this problem by resorting to the projection onto the bands obtained from a Wannier orbital model that reproduces the symmetry representations and dispersion [54, 62], and avoids the band touchings. The $6\times 6$ projector approximates this regularized projector; discrepancies between the resulting effective theories (i.e. the $6\times 6$ and regularized-projector theory) stem from the states near the band touchings, and are expected to be quantitatively minor. In particular, the main result of the nonflat bands in the $n_{B}=20$ effective theory will hold independently of the projector used.